# Differential Privacy as a Perk: Federated Learning over Multiple-Access Fading Channels with a Multi-Antenna Base Station

### v2

@@ Abstract
Federated Learning (FL) is a distributed learning paradigm that preserves privacy by eliminating the need to exchange raw data during training. In its prototypical edge instantiation with underlying wireless transmissions enabled by analog over-the-air computing (AirComp), referred to as over-the-air FL (AirFL), the inherent channel noise plays a unique role of frenemy in the sense that it degrades training due to noisy global aggregation while providing a natural source of randomness for privacy-preserving mechanisms, formally quantified by differential privacy (DP). It remains, nevertheless, challenging to effectively harness such channel impairments, as prior arts, under assumptions of either simple channel models or restricted types of loss functions, mostly considering (local) DP enhancement with a single-round or non-convergent bound on privacy loss. In this paper, we study AirFL over multiple-access fading channels with a multi-antenna base station (BS) subject to user-level DP requirements. Despite a recent study, which claimed in similar settings that artificial noise (AN) must be injected to ensure DP in general, we demonstrate, on the contrary, that DP can be gained as a perk even without employing any AN. Specifically, we derive a novel bound on DP that converges under general bounded-domain assumptions on model parameters, along with a convergence bound with general smooth and non-convex loss functions. Next, we optimize over receive beamforming and power allocations to characterize the optimal convergence-privacy trade-offs, which also reveal explicit conditions in which DP is achievable without compromising training. Finally, our theoretical findings are validated by extensive numerical results.
@@ I Introduction
Federated learning (FL) has emerged as a promising distributed learning paradigm, enabling multiple devices to collaboratively train a shared model without exposing their local data. This protocol offers significant advantages in terms of data privacy and communication efficiency over traditional centralized approaches. In modern wireless networks, the rapid proliferation of intelligent edge devices—such as smartphones and IoT sensors—has resulted in massive volumes of data being generated at the edge, beyond the reach of centralized base stations (BSs). This trend has spurred growing interest in wireless FL. As envisioned in [1], wireless FL is expected to be a key enabler of flexible, scalable, and privacy-aware intelligence acquisition from numerous data-siloed entities.
@@
Over-the-air federated learning (AirFL) represents a paradigm shift in wireless FL by enabling multiple clients to simultaneously access the uplink channel through the superposition of radio-frequency (RF) waveforms, thereby achieving scalable training that remains communication-efficient regardless of the number of participating clients [2, 3, 4]. However, the practical deployment of AirFL faces significant challenges, including channel noise and fading inherent to analog transceivers, which deteriorate learning performance by introducing errors in model aggregation [5]. Numerous prior efforts have focused on analyzing and enhancing the convergence of AirFL, the key metric for training efficacy. For instance, Zhu et al. in [4] introduced a broadband AirFL scheme using truncated channel inversion, elucidating the trade-off between signal-to-noise ratio (SNR) and the expected update-truncation ratio. Building on this, a truncated channel inversion-based AirFL system with memory mechanisms was proposed by [6], achieving optimal convergence rates under fading channels. To further reduce the communication overhead, Zhu et al. in [7] proposed a one-bit over-the-air aggregation scheme integrated with AirFL while analyzing its convergence under noisy channels. Cao et al. in [8] derived and optimized transmission powers to minimize a convergence error upper bound of AirFL. Moreover, for an edge server equipped with a multi-antenna base station (BS), Yang et al. proposed multiple-input and multiple-output (MIMO)-based AirFL and maximized the number of joint devices by joint device selection and beamforming optimization subject to the aggregation error requirement [3]. When the number of antennas is large, [9] and [10] exploited channel hardening to achieve near-optimal aggregation with reduced pilot overhead and improved scalability.
@@
On the other hand, naive privacy preserving by FL is not sufficient against emerging privacy attacks, such as membership inference attack [11, 12], and differential privacy (DP) is a formal performance metric with theoretical guarantee [13, 14, 15, 16]. In federated settings, DP is typically implemented in one of two protocols: local DP (LDP) or user-level DP. In LDP, each client perturbs its local update before transmitting it to the server, thereby safeguarding individual data points and combating inference attacks on shared data values [17, 18]. In contrast, user-level DP, which is more commonly adopted in FL, aims to protect a user’s entire contribution throughout the FL training process [19, 20]. This is typically achieved by allowing a trusted server to aggregate the parameter updates from multiple users and then applying a calibrated noise mechanism to the aggregated result [21].
@@
The most important class of mechanisms to achieve DP is the Gaussian mechanism, by which the calibrated Gaussian noise is added to the output of a function to mask the contribution of any input data. In this regard, the above-mentioned channel noise in AirFL due to wireless analog transmission is nonetheless beneficial to DP, thus becoming a frenemy in terms of privacy-preserving training. There was previous work that investigated DP employing artificial noise alongside the intrinsic channel noise in wireless FL settings. For instance, Liu and Simeone in [22] considered conditions in which DP is guaranteed without compromising training, i.e., for free. Subsequently, Hu et al. in [23] proposed an exploration into device sampling policies with replacement as a potential mechanism for augmenting the DP levels of devices by assuming that inactive devices can choose to transmit artificial noise. Also, Park and Choi in [24] considered the inherent randomness of the local gradient, which can be used to enhance the privacy analysis. However, [23, 24] only considered single-antenna BS over (quasi-) static AWGN channels, which cannot employ spatial diversity to improve training-privacy trade-offs. Although [25, 26] studied multi-antenna BS over possibly more complex channels, the analysis of convergence bounds was built upon assumptions on the loss function satisfying Polyak-Lojasiewicz conditions or being strongly convex. In addition, [27, 28] optimized the power allocation policy leveraging the channel noise, but focused on privacy loss in a single round, failing to capture the effect of mechanism composition spanning over the entire training.
@@
To gain a fundamental understanding of the role played by channel noise in DP preserving, in this paper, we consider a practical AirFL setting where multiple wireless devices (WDs) collaborate to train or fine-tune a shared model by uplink transmissions over multiple-access fading channels. We are typically interested in answering what is the optimal receive beamforming design for the multi-antenna BS to achieve convergence-privacy trade-offs with general smooth and non-convex loss functions, and whether or not there is any DP gained as a perk if we only aim for convergence improvement. Recently, Liu et al. proposed a transceiver design in a similar setup under the assumption of full client participation [26]. However, their approach suffers from several technical limitations. First, their framework provides a non-convergent, sample-level DP guarantee, which is less practical for the FL setting and offers limited privacy guarantees when the number of communication rounds becomes large. Besides, their convergence analysis is restricted to the case with strongly-convex loss functions, thus challenging its theoretical validity in modern deep learning applications. Moreover, the effect of gradient clipping is not characterized, which is an essential component for guaranteeing DP. Above all, they asserted that “privacy-for-free” is not generally achievable for AirFL in such settings, and hence (additional) artificial noise mechanisms at device side must be required, a notion that our work challenges. The main contributions of the paper are summarized as follows:
@@
  • A convergent upper bound on DP is derived under general smooth and non-convex loss functions with a bounded parameter domain assumption. It does not grow infinitely with the number of communication rounds after a burn-in period, thus offering a much tighter and more practical privacy characterization.
@@
  • We formulate and solve an optimization problem to jointly design the receive beamforming and power allocation. Our design achieves an optimal trade-off between model convergence and the user-level DP guarantee.
@@
  • Our analysis explicitly reveals that the zero-artificial-noise property always holds for general multi-antenna cases. We also reveal explicit conditions in which DP is gained as a perk in AirFL with no influence in training.
@@
  • The numerical experiments demonstrate that in practical settings, benchmarking AirFL training without privacy concerns, DP can be gained as a perk starting from a reasonably high privacy requirement and in low signal-to-noise ratio (SNR) regimes, beyond a comparable level of test accuracy.
@@ II-B Communication Model
As in [22, 26, 29], we assume that the downlink communication is ideal and noiseless, allowing each device to receive the global model $\boldsymbol{\theta}^{(t)}$ from the BS without distortion. This assumption is justified because the BS typically has far fewer resource constraints compared to the WDs for uplink transmissions. We refer to the above algorithm as AirFL-DP, which is summarized in Algorithm 1.
@@ II-C Privacy Model
As shown in Fig. 1, we consider that there is a curious third-party attacker who particularly attempts to infer information related to certain users (WDs), thus compromising user-level privacy, while the WDs and the BS are assumed to be trustful. As in [30, 31, 32], we assume a black-box setting where the attacker can only access the system’s final-released model, $\boldsymbol{\theta}^{(T)}$, after the entire FL process is complete without access to any intermediate training artifacts (e.g., gradients or parameter updates).
@@ III-B User-level Privacy Guarantee
The remaining term in the RHS of (16), with $\boldsymbol{\varsigma}^{(t)}$, is aimed for further bounding the shifted Rényi divergence leveraging the following lemma [30, Lemma 20].
@@
Proof: The detailed proof is provided in Appendix B. $\blacksquare$
@@
where $\kappa_{\max}=\max_{t}\eta L\mathord{\left/\right.\kern-1.2pt}(rn)\sum_{i\in\mathcal{I}^{(t)}}(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}s_{i}^{(t)}$.
Proof: We summarize the sketch of proof as follows. This proof proceeds by a recursion on the iteration index $t$, mainly involving the following key steps:
  1. 1.
    Decomposition: At each iteration $t$, we conceptually split the AirFL-DP update procedure into two components: (i) a noisy update function and (ii) a shift-reduction mechanism that incorporates the remaining Gaussian noise.
  2. 2.
    Base case ($t=\tau$, for any $\tau\in\{0,1,\ldots,T-1\}$): We leverage Assumption 2.
  3. 3.
    Recursive step: As for the recursive step, we first bound the privacy cost of the noisy update component using Lemma III.3. Next, we analyze the shift-reduction mechanism using Lemma III.2, which involves introducing auxiliary shift variables to control the shift parameter.
  4. 4.
    Trace RDP loss: The overall upper bound on RDP is derived by repeatedly applying the recursive step from $T$ to $\tau$ and appropriately setting the values of the auxiliary shift variables.
The detailed proof is provided in Appendix C. $\blacksquare$
As shown in Fig. 2, a key feature of our analysis result is that the privacy loss can converge to a constant value irrespective of the number of communication rounds. In other words, the AirFL-DP algorithm will stop accumulating privacy loss after a sufficient number of iterations. This marks a significant improvement compared to previous analyses, where the privacy bound typically grows unboundedly with $T$ [22, Lemma 3], [26, Corollary 1]. Furthermore, our analysis gracefully recovers the linear bound when the domain diameter $D$ approaches infinity.
Another important insight gained from our analysis is the critical role played by the wireless channel impairments. Our proposition demonstrates that the DP cost is paid by inherent channel noise for free. It suggests that an arbitrary level $\epsilon$ is achievable by controlling the receive beamforming vector $\boldsymbol{w}^{(t)}$, power scaling factor $s_{i}^{(t)}$, and the clipping threshold $c$ in conditions of fading channel $\boldsymbol{h}_{i}^{(t)}$ and channel noise of power $\sigma^{2}$, $i\in\mathcal{I}^{(t)}$.
Figure 2: An illustration of the RDP privacy loss $\epsilon^{\prime}$ as a function of the total number of communication rounds $T$. For ease of exposition, this plot sets a constant $\phi_{t}=\phi$ for all $t$. The full analysis and parameter settings used in our experiments are specified in Section V-A.
@@
Proof: The detailed proof is provided in Appendix D. $\blacksquare$
@@ IV-A Convergence Analysis
Assumption 4 (Bounded SGD Variance and Norm)
@@
Our convergence analysis builds on a recent result in [5], which addresses the convergence of our algorithm in a simplified setting. Specifically, they considered additive white Gaussian noise (AWGN) channel (i.e., $\boldsymbol{h}_{i}^{(t)}=\boldsymbol{1}$, for all $i\in[n]$ and $t$), full client participation ($r=1$), and a single-antenna base station ($m=1$). Their analysis relies on the assumption that the variance of estimation error $\boldsymbol{n}_{\text{est}}^{(t)}=\hat{\boldsymbol{\Delta}}^{(t)}-\sum_{i\in\mathcal{I}}\bar{\boldsymbol{\Delta}}_{i}^{(t)}$ can be expressed as
@@
Proof: The proof is based on Lemma IV.1 and is detailed in Appendix E. $\blacksquare$

### v3

@@ Abstract
Federated Learning (FL) is a distributed learning paradigm that preserves privacy by eliminating the need to exchange raw data during training. In its prototypical edge instantiation with underlying wireless transmissions enabled by analog over-the-air computing (AirComp), referred to as over-the-air FL (AirFL), the inherent channel noise plays a unique role of frenemy in the sense that it degrades training due to noisy global aggregation while providing a natural source of randomness for privacy-preserving mechanisms, formally quantified by differential privacy (DP). It remains, nevertheless, challenging to effectively harness such channel impairments, as prior arts, under assumptions of either simple channel models or restricted types of loss functions, mostly consider (local) DP enhancement with a single-round or non-convergent bound on privacy loss. In this paper, we study AirFL over multiple-access fading channels with a multi-antenna base station (BS) subject to user-level DP requirements. Despite a recent study, which claimed in similar settings that artificial noise (AN) must be injected to ensure DP in general, we demonstrate, on the contrary, that DP can be gained as a perk even without employing any AN. Specifically, we derive a novel bound on DP that converges under general bounded-domain assumptions on model parameters, along with a convergence bound with general smooth and non-convex loss functions. Next, we optimize over receive beamforming and power allocations to characterize the optimal convergence-privacy trade-offs, which also reveal explicit conditions in which DP is achievable without compromising training. Finally, our theoretical findings are validated by extensive numerical results.
@@ I Introduction
Federated learning (FL) has emerged as a promising distributed learning paradigm, enabling multiple devices to collaboratively train a shared model without exposing their local data. This protocol offers significant advantages in terms of data privacy and communication efficiency over traditional centralized approaches. In modern wireless networks, the rapid proliferation of intelligent edge devices—such as smartphones and IoT sensors—has resulted in massive volumes of data being generated at the edge, beyond the reach of centralized base stations (BSs). This trend has spurred growing interest in wireless FL. As envisioned in [36], wireless FL is expected to be a key enabler for flexible, scalable, and privacy-aware intelligence acquisition from numerous data-siloed entities.
@@
Over-the-air federated learning (AirFL) represents a paradigm shift in wireless FL by enabling multiple clients to simultaneously access the uplink channel through the superposition of radio-frequency (RF) waveforms, thereby achieving scalable training that remains communication-efficient regardless of the number of participating clients [25, 38, 41]. However, the practical deployment of AirFL encounters significant challenges, including channel noise and fading inherent to analog transceivers, which deteriorate learning performance by introducing errors in model aggregation [33]. Numerous prior efforts have focused on analyzing and enhancing the convergence of AirFL, the key metric for training efficacy. For instance, Zhu et al. in [41] introduced a broadband AirFL scheme using truncated channel inversion, elucidating the trade-off between signal-to-noise ratio (SNR) and the expected update-truncation ratio. Building on this, a truncated channel inversion-based AirFL system with memory mechanisms was proposed by [34], achieving optimal convergence rates under fading channels. To further reduce communication overhead, Zhu et al. in [39] proposed a one-bit over-the-air aggregation scheme integrated into AirFL while analyzing its convergence under noisy channels. Cao et al. in [4] derived and optimized transmission power to minimize the convergence error upper bound of AirFL. Moreover, for an edge server equipped with a multi-antenna base station (BS), Yang et al. proposed multiple-input and multiple-output (MIMO)-based AirFL and maximized the number of joint devices by joint device selection and beamforming optimization subject to aggregation error requirements [38]. When the number of antennas is large, [32] and [6] exploited channel hardening to achieve near-optimal aggregation with reduced pilot overhead and improved scalability.
@@
On the other hand, naive privacy preserving by FL is not sufficient against emerging privacy attacks, such as membership inference attack [27, 24], and differential privacy (DP) is a formal performance metric with theoretical guarantee [9, 1, 22]. In federated settings, DP is typically implemented in one of two protocols: local DP (LDP) or user-level DP. In LDP, each client perturbs its local update before transmitting it to the server, thereby safeguarding individual data points and combating inference attacks on shared data values [29, 5]. By contrast, user-level DP, which is more commonly adopted in FL, aims for protecting one user’s entire contribution throughout the FL training process [21, 11]. This is typically achieved by allowing a trusted server to aggregate the parameter updates from multiple users and then applying a calibrated noise mechanism to the aggregated result [31].
@@
The most important class of mechanisms to achieve DP is the Gaussian mechanism, by which the calibrated Gaussian noise is added to the output of a function to mask the contribution of any input data. In this regard, the above-mentioned channel noise in AirFL due to wireless analog transmission is nonetheless beneficial to DP, thus becoming a frenemy in terms of privacy-preserving training. There was previous work that investigated DP employing artificial noise alongside intrinsic channel noise in wireless FL settings. For instance, Liu and Simeone in [16] considered conditions in which DP is guaranteed without compromising training, i.e., for free. Subsequently, Hu et al. in [13] proposed exploration into device sampling with replacement as a potential mechanism for augmenting the DP levels of devices by assuming that inactive devices can choose to transmit artificial noise. Also, Park and Choi in [26] considered the inherent randomness of the local gradient, which can be used to enhance the privacy analysis. However, [13, 26] only considered single-antenna BS over (quasi-) static AWGN channels, which cannot employ spatial diversity to improve training-privacy trade-offs. Although [17, 18] studied multi-antenna BS over possibly more complex channels, the analysis of convergence bounds was built upon assumptions on the loss function satisfying Polyak-Lojasiewicz (PL) conditions or being strongly convex. In addition, [14, 37] optimized the power allocation policy leveraging the channel noise, but focused on privacy loss in a single round, failing to capture the effect of mechanism composition spanning over the entire training.
@@
To gain a fundamental understanding of the role played by channel noise in DP preserving, in this paper, we consider a practical AirFL setting where multiple wireless devices (WDs) collaborate to train or fine-tune a shared model by uplink transmissions over multiple-access fading channels. We are typically interested in answering what is the optimal receive beamforming design for the multi-antenna BS to achieve convergence-privacy trade-offs with general smooth and non-convex loss functions, and whether or not there is any DP gained as a perk if we only aim for convergence improvement. Recently, Liu et al. proposed a transceiver design in a similar setup under the assumption of full client participation [18]. However, their approach suffers from several technical limitations. First, their framework provides a non-convergent and sample-level DP guarantee, which is less practical for the FL setting, and offers limited privacy guarantees when the number of communication rounds becomes large. Besides, their convergence analysis is restricted to the case with strongly-convex loss functions, thus challenging its theoretical validity in modern deep learning applications. Moreover, the effect of gradient clipping is not formally characterized, which is an essential component for guaranteeing DP. Above all, they asserted that “privacy-for-free” is not generally achievable in such settings, and hence (additional) artificial noise mechanisms at the device side is inevitable in analysis, a notion that our work challenges. Table I compares our work with existing literature in terms of problem problem problem, assumptions, formal capture of gradient clipping, achievable DP, and training convergence. The main contributions of the paper are summarized as follows:
@@
  • A convergent upper bound on DP is derived under general smooth and non-convex loss functions with a bounded parameter domain assumption. It does not grow infinitely with the number of communication rounds after a burn-in period, thus offering much tighter and more practical privacy characterization.
@@
  • We formulate an optimization problem to jointly design the receive beamforming and power allocation. Our proposed solution achieves an optimal trade-off between model convergence and the user-level DP guarantee.
@@
  • Our analysis explicitly reveals that zero-artificial-noise property is always possible for general multi-antenna cases. We also identify explicit conditions in which DP is gained as a perk in AirFL settings with no compromise to training.
@@
  • The numerical experiments demonstrate that benchmarking AirFL training without privacy concerns, AirFL can gain DP as a perk with reasonably high privacy requirement and in low signal-to-noise ratio (SNR) regimes, while achieving a comparable level of test accuracy.
TABLE I: Comparison of the privacy guarantees and the assumptions required by different works. $\ddagger$}” indicates that additional parameter estimation (e.g., the norm and covariance of the random gradient) is required. $\dagger$}” claims that relying solely on channel noise cannot meet high privacy requirements in multi-user SIMO settings.
Reference Channel Model Threat Model Loss Function Gradient Clipping? Convergent Privacy? Convergence?
Koda et al. [14] multi-user SISO sample-level DP Lipschitz $\times$ $\times$ $\times$
time-invariant channel honest but curious BS
Liu et al. [16] multi-user SISO sample-level DP smooth, Lipschitz, PL condition $\times$ $\times$ $\surd$
block flat-fading channel honest but curious BS
Park et al. [26] multi-user SISO sample-level DP smooth, Lipschitz $\times$ $\times$ $\surd$
block flat-fading channel honest but curious BS
Liu et al. [18] multi-user SIMO sample-level DP smooth, Lipschitz, strongly convex $\times$ $\times$ $\surd$
time-invariant channel honest but curious BS
Ours multi-user SIMO user-level DP smooth $\surd$ $\surd$ $\surd$
block flat-fading channel curious third-party attacker
@@ II-B Communication Model
As in [16, 18], we assume that the downlink communication is ideal and noiseless, allowing each device to receive the global model $\boldsymbol{\theta}^{(t)}$ from the BS without distortion. This assumption is justified because the BS typically has far fewer resource constraints compared to the WDs for uplink transmissions. We refer to the above algorithm as AirFL-DP, which is summarized in Algorithm 1.
@@ II-C Privacy Model
We consider a curious third-party attacker who particularly attempts to infer information about individual users (WDs), thereby compromising user-level privacy, while both the WDs and the BS are assumed to be trustworthy. Under this setting, the attacker can only access the system’s final-released model, $\boldsymbol{\theta}^{(T)}$, upon completion of the entire FL training without access to any intermediate model artifacts (e.g., gradients or parameter updates) [10, 2, 15]. Note that if the BS is assumed to be a “curious but honest” adversary [16, 18], it would know which WDs participate in each round through its role in coordinating uplink transmissions and scheduling, making the notion of user-level DP invalid in such a setting.
@@ III-B User-level Privacy Guarantee
The remaining term in the RHS of (16), with $\boldsymbol{\varsigma}^{(t)}$, is aimed for further bounding the shifted Rényi divergence leveraging the following lemma [10, Lemma 20].
@@
Proof: The detailed proof is provided in Appendix A. $\blacksquare$
@@
where $\kappa_{\max}=\max_{t}\eta L\sum_{i\in\mathcal{I}^{(t)}}(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}s_{i}^{(t)}\mathord{\left/\right.\kern-1.2pt}(rn)$.
Proof: The detailed proof is provided in Appendix B. $\blacksquare$
As shown in Fig. 2, a key feature of our analysis result is that the privacy loss can converge to a constant value irrespective of the number of communication rounds. In other words, the AirFL-DP algorithm will stop accumulating privacy loss after a sufficient number of iterations. This marks a significant improvement compared to previous analyses, where the privacy bound typically grows unboundedly with $T$ [16, Lemma 3], [18, Corollary 1]. Furthermore, our analysis gracefully recovers the linear bound when the domain diameter $D$ approaches infinity.
Another important insight gained from our analysis is the critical role played by the wireless channel impairments. Our proposition demonstrates that the DP cost is paid by inherent channel noise for free. It suggests that an arbitrary level $\epsilon^{\prime}$ is achievable by controlling the receive beamforming vector $\boldsymbol{w}^{(t)}$, power scaling factor $s_{i}^{(t)}$, and the clipping threshold $c$ in conditions of fading channel $\boldsymbol{h}_{i}^{(t)}$ and channel noise of power $\sigma^{2}$, $i\in\mathcal{I}^{(t)}$.
Figure 2: An illustration of the RDP privacy loss $\epsilon^{\prime}$ as a function of the total number of communication rounds $T$, compared with [18]. For clarity, we set $\phi_{t}=\phi$ for all $t$ in this plot. The complete analysis and parameter settings used in our experiments are specified in Sec. V-A.
@@
Proof: The detailed proof is provided in Appendix C. $\blacksquare$
@@ IV-A Convergence Analysis
Assumption 4 (Bounded SGD Variance)
@@
Our convergence analysis builds on a recent result in [33], which addresses the convergence of our algorithm in a simplified setting. Specifically, they considered an additive white Gaussian noise (AWGN) channel (i.e., $\boldsymbol{h}_{i}^{(t)}=\boldsymbol{1}$, for all $i\in[n]$ and $t$), full client participation ($r=1$), and a single-antenna base station ($m=1$). Their analysis relies on the assumption that the variance of estimation error $\boldsymbol{n}_{\text{est}}^{(t)}=\hat{\boldsymbol{\Delta}}^{(t)}-\sum_{i\in\mathcal{I}}\bar{\boldsymbol{\Delta}}_{i}^{(t)}$ can be expressed as
@@
Proof: Replace $\mathbb{E}\left[\|\boldsymbol{n}_{\text{est}}^{(t)}\|^{2}\right]=dv^{(t)}$ in Lemma IV.1 with the following upper bound
$\mathbb{E}\left[\|\boldsymbol{n}_{\text{est}}^{(t)}\|^{2}\right]\leq rn\Lambda^{(t)}+d\|\boldsymbol{w}^{(t)}\|^{2}\sigma^{2},$ (30)
and combine the fact that the random selection of $\mathcal{I}^{(t)}$ is uniform, with $|\mathcal{I}^{(t)}|=rn$, yielding the desired result. $\blacksquare$